# What is the modular Fourier transform?

### v2

@@ 7.1. Case $p\nmid N$
Note that solutions to $x^{N}-1=0$ in $F_{p}$ are just roots of unity in a finite field. Since $x^{p-1}-1=0$ in any field of order $p$, we require that $N|p-1$ to have solutions. Further, when there are solutions there are $gcd(N,p-1)$ of them.
@@ 7.1.1. Examples
Note that we may use the Chinese remainder theorem to obtain an isomorphism to the decomposition into summands: [DFTCG]
$F_{p}\left[x\right]/(x^{N}-1)\cong\prod_{d|N,i}F_{p}\left[x\right]/(P_{i}(x))$
The map is just given by $f\mapsto f\mod P_{i}(x)$ for each of the $g$ $P_{i}$’s occuring in the factorization of the cyclotomic polynomials. The inverse map is given by multiplying the factors in the variable $x$, and taking the residue $\mod x^{N}-1$.
We may want to find an extension $F_{q}$ of $F_{p}$ where $q=p^{\nu}$ which splits $x^{N}-1$. This is always possible, since we just need to find a primitive $N^{th}$ root, which occurs when $N|q-1$. In this case, we have
@@ 7.2. Case $p\mid N$
Writing $N=p^{s}l$ where $p\nmid l$ we have
$x^{N}-1=(x^{l}-1)^{p^{s}}$
so it follows that $r\leq p^{s}$ (but now it’s possible that $r>1$). If $r>1$, then the corresponding representation $F\left[T\right]/q(T)^{r}$ is indecomposable and not irreducible, where $q$ is an irreducible factor of $x^{l}-1$ over $F$. The irreducible representations occur precisely when $r=1$.
@@
$T^{p^{s}l}-1=(T^{l}-1)^{p^{s}}=0$
Thus $T^{l}-1$ is an intertwining operator which is not invertible, so by Schur’s lemma it is equal to zero.
In this case, we may still factor $x^{l}-1$ as before, we just obtain multiplicities on all factors, so the representations are now indecomposable rather than irreducible. Thus, we have the isomorphism
$F_{p}\left[x\right]/(x^{N}-1)\cong F_{p}\left[x\right]/(P_{i}(x)^{p^{s}})$
@@ 8. The Symmetric Group over $\mathbb{F}_{p}$
We first need to compute the compute the dimension numbers for the simple modules $D^{\mu}$, which can be obtained by computing the rank of the Gram matrix over $\mathbb{F}_{p}$ for the polytabloid basis [MRTCC] [DMRSG].
We need to understand the structure of the blocks in terms of simple modules. The blocks have Jordan-Hölder composition series with simple modules $D^{\mu}$ for $p$-regular partitions $\lambda$ occurring as simple factors. The number of occurrences of $D^{\mu}$ in the series for $S^{\lambda}$ is usually denoted $d^{\lambda\mu}$, and the matrix $D$ is called the decomposition matrix. These are computed beginning on pg. 98 in [GJ78] [GJ88]. The top composition factor is $D^{\lambda}=S^{\lambda}/(S^{\lambda}\cap(S^{\lambda})^{\perp})$.
@@
The idempotents are computed by Murphy [GM83] as sums involving the Jucys-Murphy elements, and equivalence classes of partitions based on their $p$-core. The $p$-core of a partition is obtained by removing as many rim $p$-hooks from the edge of a partition as possible. The number removed is the weight. $p$-cores $\gamma$ parameterize the blocks of the symmetric group. [MONC]
Once we have computed the idempotents, we can write $F_{p}\left[S_{N}\right]e_{i}=\text{span}_{F_{p}}\{\sigma e_{i}|\sigma\in S_{N}\}$ to find a basis for the block. Putting these bases together, we obtain a second basis for the symmetric group algebra different from the basis $\{\sigma|\sigma\in S_{N}\}$. The change of basis matrix is the modular Fourier transform. [DFTSG]
@@ 8.1.1. $p=2$, $n=3$
As expected, we obtain a matrix of order 4 despite the fact that $2\mid 6$.
@@ 8.1.2. $p=3$, $n=4$
print((modular_fourier_transform(3,4)^488488).str())
[1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1 0]
[0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 1]
@@ 8.2. $p=7$, $n=4$
9. Applications
When $p>n$, we are no longer in the modular case but can still compute these matrices. When $p$ is very large, one can use the matrix to encode a vector by multiplication, and decode by multiplying by the inverse. This would appear to be a nascent symmetric key cryptosystem with key $p$. We make no claims to its security or efficiency.
In quantum computing, the technique of Fourier sampling is used prominently in Shor’s algorithm for the abelian group $\mathbb{Z}/N\mathbb{Z}$. In graph isomorphism, the hidden subgroup is the automorphism group of a graph, which is a subgroup of the symmetric group. Strong Fourier sampling fails for the symmetric group over $\mathbb{C}$, but to our knowledge has not been explored over finite fields. [SFS]
Appendix A Computing Central Orthogonal Idempotents
We provide a direct implementation of [GM83] to compute central primitive orthogonal idempotents [IDEM], but it is significantly slower than the built in method to compute idempotents.
timeit.timeit("central_orthogonal_idempotents(3,5)", number=1, globals=globals())
14.356356750002305
timeit.timeit("SymmetricGroupAlgebra(GF(3),5) \
.central_orthogonal_idempotents()", number=1, globals=globals())
0.02521254200109979

### v3

@@ 7.1. Case $p\nmid N$
Note that solutions to $x^{N}-1=0$ in $F_{p}$ are just roots of unity in a finite field. Since $x^{p-1}-1=0$ in any field of order $p$, we require that $N|(p-1)$ to have solutions. Further, when there are solutions there are $gcd(N,p-1)$ of them.
@@ 7.1.1. Examples
Note that we may use the Chinese remainder theorem to obtain an isomorphism to the decomposition into summands: [DFT]
$F_{p}\left[x\right]/(x^{N}-1)\cong\bigoplus_{d|N,i}F_{p}\left[x\right]/(P_{i}(x))$
The map is just given by $f\mapsto f\mod P_{i}(x)$ for each of the $g$ $P_{i}$’s occuring in the factorization of the cyclotomic polynomials. The inverse map is given by multiplying the factors in the variable $x$, and taking the residue modulo $x^{N}-1$.
We may want to find an extension $F_{q}$ of $F_{p}$ where $q=p^{\nu}$ that splits $x^{N}-1$. This is always possible, since we just need to find a primitive $N^{th}$ root, which occurs when $N|(q-1)$. In this case, we have
@@ 7.2. Case $p\mid N$
Writing $N=p^{s}\ell$ where $p\nmid\ell$ we have
$x^{N}-1=(x^{\ell}-1)^{p^{s}}$
so it follows that $r\leq p^{s}$ (but now it is possible that $r>1$). If $r>1$, then the corresponding representation $F\left[T\right]/q(T)^{r}$ is indecomposable and not irreducible, where $q$ is an irreducible factor of $x^{\ell}-1$ over $F$. The irreducible representations occur precisely when $r=1$.
@@
$T^{p^{s}\ell}-1=(T^{\ell}-1)^{p^{s}}=0$
Thus $T^{\ell}-1$ is an intertwining operator which is not invertible, so by Schur’s lemma it is equal to zero.
In this case, we may still factor $x^{\ell}-1$ as before, we just obtain multiplicities on all factors, so the representations are now indecomposable rather than irreducible. Thus, we have the isomorphism
$F_{p}\left[x\right]/(x^{N}-1)\cong F_{p}\left[x\right]/(P_{i}(x)^{p^{s}}).$
@@ 8. The Symmetric Group over $\mathbb{F}_{p}$
Though we do not make use of them, the simple modules $D^{\lambda}=S^{\lambda}/\left(S^{\lambda}\cap(S^{\lambda})^{\perp}\right)$ occur for each $p$-regular partition $\lambda$, and form a complete traversal. They are computed in SageMath 10.4+ [SAGE] for small $n$, and much more extensive tables are found in [GJ78, pg. 98]; see also [GJ88].
The idempotents are computed by Murphy [GM83] as sums involving the Jucys–Murphy elements, and equivalence classes of partitions based on their $p$-core. The $p$-core of a partition is obtained by removing as many rim $p$-hooks from the edge of a partition as possible. The number removed is the weight. $p$-cores $\gamma$ parameterize the blocks of the symmetric group. [MW]
We include an explicit computation of the idempotents in [DFT], and they can be computed much more quickly using Sage via
p=3; n=4;
SGA = SymmetricGroupAlgebra(GF(p),n)
print(SGA.central_orthogonal_idempotents())
Once we have computed the idempotents, we can write $F_{p}\left[S_{N}\right]e_{i}=\text{span}_{F_{p}}\{\sigma e_{i}|\sigma\in S_{N}\}$ to find a basis for the block. Putting these bases together, we obtain a second basis for the symmetric group algebra different from the basis $\{\sigma|\sigma\in S_{N}\}$. The change of basis matrix is the modular Fourier transform [DFT].
@@ 8.1.1. $p=2$, $n=3$
We obtain a matrix of order 4 despite the fact that $2\mid 6$.
@@ 8.1.2. $p=3$, $n=4$
print((modular_fourier_transform(3,4)^488488) == identity_matrix(factorial(4)))
True
@@ 8.2. $p=7$, $n=4$
9. Unitarity for the cyclic group
The discrete Fourier transform for the cyclic group over the complex numbers is not unitary when given by $A=\left(\omega^{ij}\right)_{0\leq i,j<N}$ where $\omega=\exp{2\pi i/N}$. $A^{4}=N^{2}I_{N}$. To make it unitary, we multiply by a factor of $\frac{1}{\sqrt{N}}$ to obtain $U=\frac{1}{\sqrt{N}}A$, and $U^{4}=I_{N}$. This means that the eigenvalues of $U$ satisfy $\lambda^{4}=1$, and so are fourth roots of unity $\{-1,+1,-i,+i\}$.
Over finite fields, the situation is slightly more delicate but is in complete analogy to the case over the complex numbers. First we need a notion of conjugation appropriate for finite fields. Suppose we are working over a finite field $F_{q}$. If $q$ is square, then $q=(q^{\prime})^{2}$ and if not embed $F_{q}\rightarrow F_{q^{2}}$, so without loss of generality we can work over $F_{q^{2}}$. Then define $\alpha:x\rightarrow x^{q}$ to be an order two involution, the $r^{th}$ power of Frobenius if $q=p^{r}$. Then an operator is unitary if the inner product $\langle x,y\rangle=\sum_{i}x_{i}^{q}y_{i}$ on an $F_{q^{2}}$ vector space is preserved. In other words, if $UU^{*}=I_{N}$ where $U^{*}=(U^{\alpha})^{T}$.
When $p\nmid N$, we have an $N^{th}$ root of unity $\alpha$, so the unnormalized DFT is $A_{ij}=\alpha^{ij}$. $\sqrt{N}$ may not be in $F_{q^{2}}$. If not, form the quadratic extension $F_{q^{2}}[x]/(x^{2}-N)\cong F_{q^{4}}$ which surely contains $\sqrt{N}$.
$\left(AA^{*}\right)_{ik}=\sum_{j=0}^{n-1}\alpha^{j(i+qk)}=N\delta_{i,-qk}$
where we use a geometric sum to simplify. Normalizing, $UU^{*}=\delta_{i,-qk}$. Thus $U$ is unitary iff $q\equiv-1(\mod N)$. To compute the order, $(U^{2})_{ik}=\sum_{j=0}^{N-1}\alpha^{j(i+k)}=\delta_{i,-k}$ by another geometric sum, and $(U^{4})_{ik}=(U^{2}U^{2})_{ik}=\sum_{j=0}^{N-1}\delta_{i,-j}\delta_{j,-k}=\delta_{ik}$, thus $U$ has order 4. Since the unitary DFT is order 4, the eigenvalues are four fourth roots of unity when working over a splitting field $F_{\tilde{q}}$ of the characteristic polynomial. If $a$ is a primitive element of $F_{\tilde{q}}$, then the eigenvalues are {±$\{\pm 1,\pm a^{(\tilde{q}-1)/4}\}$.
The eigenvectors of the DFT of the cyclic group over the complex numbers are computed in [PM80]. Note that Schur’s matrix is a cyclic permutation $(12\,\ldots\,N)$ applied to the rows and columns of the DFT matrix, so they are conjugate.
10. Unitarity for the symmetric group
The Fourier transform described in §3 is not unitary. In order make the transform unitary, we need to include factors involving square roots. We’ll start by working over the complex numbers, and then consider more economical subfields. Consider
$\hat{f}(\tau)=\sum_{g\in G}\sqrt{\frac{d_{\tau}}{|G|}}f(g)\tau(g)$
where every $\tau(g)\in U(n)$ is now a unitary matrix, and $d_{\tau}$ is the dimension of the representation $\tau$. The overall DFT is now unitary with these two modifications [RB]. Note that every representation of a finite group, or even a compact Lie group, may be made unitary by Weyl’s unitary trick. Recall the invariant inner product $\langle\cdot,\cdot\rangle_{G}=\frac{1}{|G|}\sum_{g\in G}\langle\rho(g)x,\rho(g)y\rangle$. By using the change-of-basis matrix associated to this inner product, we can conjugate $\rho$ to a unitary representation $\tau$. There are different versions of the unitary trick, but we will use the following. Let
[…]
and $P=Q^{2}$, $Q$ is a principal square root of $P$, and $dg$ is the Haar measure. $Q$ is the change-of-basis matrix required, and $\tau(g)=Q^{-1}\rho(g)Q$ is unitary. For finite groups, $P=\frac{1}{|G|}\sum_{g\in G}\rho(g)\rho(g)^{*}$. $Q$ can be computed by diagonalizing the positive symmetric matrix $P$. Let $P=MDM^{-1}$, then $Q=MD^{1/2}M^{-1}$.
10.1. Over number fields
Before computing the eigenvalues, it is useful to work over smaller subfields of $\mathbb{C}$ which only contain the square roots necessary to compute the unitary DFT. First consider the symmetric group algebra over the rationals, $\mathbb{Q}[S_{n}]$.
The only square roots appearing are in the “unitary factors” $\sqrt{\frac{d_{\tau}}{|G|}}$, and $D^{1/2}$. Thus, we can form successive quadratic extensions $K=\mathbb{Q}(\sqrt{d_{\tau}},\sqrt{|G|},\sqrt{d_{i}})_{\tau,i}$ for every representation $\tau$ and corresponding diagonal matrix $D$ with entries $d_{i}$ for $1\leq i\leq d_{\tau}$. $K$ is a finite degree extension of $\mathbb{Q}$, so a number field. Each $\tau:G\rightarrow GL(d_{\tau},K)$ is a unitary representation defined over $K$, and the overall DFT is unitary.
10.2. Eigenvalues for $n=3$
While we can compute the eigenvalues over $\overline{\mathbb{Q}}$, it is difficult to give explicit formulas for them except in small examples. Consider the symmetric group $S_{3}$. We must adjoin two square roots, so $K=\mathbb{Q}(\sqrt{2},\sqrt{3})$ and $[K:\mathbb{Q}]=4$. The unitary DFT matrix $U_{DFT}$ is
$\begin{bmatrix}-1/(6\sqrt{6})&-1/(6\sqrt{6})&-1/(6\sqrt{6})&-1/(6\sqrt{6})&-1/(6\sqrt{6})&-1/(6\sqrt{6})\\ -1/(3\sqrt{3})&0&-1/2&1/(6\sqrt{3})&1/(6\sqrt{3})&1/2\\ 0&-1/(3\sqrt{3})&1/(6\sqrt{3})&-1/2&1/2&1/(6\sqrt{3})\\ 0&-1/(3\sqrt{3})&1/(6\sqrt{3})&1/2&-1/2&1/(6\sqrt{3})\\ -1/(3\sqrt{3})&0&1/2&1/(6\sqrt{3})&1/(6\sqrt{3})&-1/2\\ -1/(6\sqrt{6})&1/(6\sqrt{6})&1/(6\sqrt{6})&-1/(6\sqrt{6})&-1/(6\sqrt{6})&1/(6\sqrt{6})\\ \end{bmatrix}$
The characteristic polynomial $f(x)\in K[x]$ and is given by
$f(x)=x^{6}+(-1/(3\sqrt{3})-1/2)x^{5}+((-1/(6\sqrt{2})+1/6)\sqrt{3}-1/(3\sqrt{2})-1/3)x^{4}\$
$+((1/(3\sqrt{2})+1/3)\sqrt{3}+1/(2\sqrt{2}))x^{3}+((-1/(6\sqrt{2})+1/6)\sqrt{3}-1/(3\sqrt{2})-1/3)x^{2}+(-1/(3\sqrt{3})-1/2)x+1$
$f$ is monic and palindromic. The eigenvalues (roots) lie on the unit circle since $U_{DFT}$ is unitary. Using Sage, we compute the splitting field $L$ to be of absolute degree 192, i.e. $[L:\mathbb{Q}]=192$. The defining polynomial has large integer coefficients. Therefore the relative degree is $[L:K]=[L:\mathbb{Q}]/[K:\mathbb{Q}]=192/4=48$. Over $\overline{\mathbb{Q}}$, the eigenvalues are
$\begin{matrix}[-0.9865699009578278?-0.1633396171296542?i,&-0.9865699009578278?+0.1633396171296542?i,\\ -0.5437444089082021?-0.8392508670124029?i,&-0.5437444089082021?+0.8392508670124029?i\\ 0.9916391752712170?-0.1290416447020912?i,&0.9916391752712170?+0.1290416447020912?i]\\ \\ \end{matrix}$
with minimal polynomial $g(x)\in\mathbb{Q}[x]$
$g(x)=x^{24}+2x^{23}-(1/2)x^{22}-(17/6)x^{21}-(215/144)x^{20}-(1/6)x^{19}+(509/216)x^{18}+(11/3)x^{17}-(35/1296)x^{16}$
$-(139/54)x^{15}-(1483/648)x^{14}+(7/18)x^{13}+(155/54)x^{12}+(7/18)x^{11}-(1483/648)x^{10}-(139/54)x^{9}$
$-(35/1296)x^{8}+(11/3)x^{7}+(509/216)x^{6}-(1/6)x^{5}-(215/144)x^{4}-(17/6)x^{3}-(1/2)x^{2}+2x+1$
The eigenvalues cannot be roots of unity. If they were, $g(x)$ would have factors which are cyclotomic polynomials which have all integer coefficients, and $g(x)$ has rational coefficients that are not integers.
11. Galois group of characteristic polynomial for $n=3$
Though it is difficult or impossible to compute the Galois group using Sage, we can compute it by hand. In this case, $|\text{Gal}(L/K)|=[L:K]=48$. Recall that the Galois group of a polynomial is a transitive subgroup of a permutation group. $\text{Gal}(L/K)\subset S_{6}$, and while there is only one isomorphism class of subgroups of order 48 in $S_{6}$, namely $S_{2}\times S_{4}$, there is only one such transitive permutation subgroup [DM], namely the wreath product $S_{2}\wr S_{3}$. This partitions the roots into three pairs, and permutes the pairs with $S_{3}$ while internally swapping the pairs with $S_{2}$. For a sextic, only the transitive subgroups of order 48 and 72 are solvable, so it is possible to find expressions for the eigenvalues. [TH]